A manufacturer produces two models of the same toy; Model A and Model B. Model A takes 4 hours to produce and costs $8/each. Model B takes 3 hours to produce and costs $7/each. If the manufacturer allots a total of 5800 hours and $12,600 for production each week, how many of each model will be produced?
Thanks in advance for any assistance!

Answer by mananth(16946)

Let the number of toys of Model A produced be x
Let the number of toys of Model B produced be y
..
time equation
4x+3y=5800
Cost equation
8x+7y=12600
..
Multiply the first equation by 2 and subtract the second to eliminate x.
8x+6y-8x-7y=11600-12600
-y=-1000
y= 1000 Number of toys of Model B
plug the value of y in the first equation
4x+3*1000= 5800
4x=5800-3000
x= 2800/ 4= 700
Number of toys of Model A= 700